OpenStudy (anonymous):

is 3x^2+2x-7 factorable

OpenStudy (lgbasallote):

do you think it is?

OpenStudy (anonymous):

well, try it out (3x ) (x )

OpenStudy (anonymous):

no but i cant get x ive asked this qu like 3 times and no help

OpenStudy (lgbasallote):

well yep it's not factorable heh

OpenStudy (anonymous):

i dint think soo

hero (hero):

Yes, it's factorable

OpenStudy (anonymous):

then what is it ima sooo lost!!

hero (hero):

Any trinomial is factorable. The question is, will the roots be integer, real, or complex?

OpenStudy (anonymous):

i know how to use it

hero (hero):

That's not exactly the way I would go about it. But, you probably don't want to know my methods. It might go beyond the means

OpenStudy (australopithecus):

and the test is where ax^(2) +bx + c \[\Delta x = \sqrt{b^{2}-4ac}\] when \[\Delta x = 0\] there is only one root \[\Delta x = \sqrt{Positive Number}\] there are two roots \[\Delta x = \sqrt{NegativeNumber}\]No real roots exist

OpenStudy (anonymous):

.... can i have the answer pluzz

OpenStudy (anonymous):

ok i know that stuff and thanks it means a lot :) but could u just tell me what x is

OpenStudy (anonymous):

ok i know this stuff

OpenStudy (anonymous):

no i didnt cuz u said so much other stufff sorry not trying to bee a buttface here

OpenStudy (anonymous):

dude i know how to do it i just gotta mass head ace and stuff and wanna find x thats all not some lesson cuz i know how to get it i just wanted an answer :,(

OpenStudy (anonymous):

ugh!! never mind

hero (hero):

@Australopithecus, you have not properly factored the trinomial.